A 27.2 kg dog is running northward at 2.64 m/s , while a 5.85 kg cat is running eastward at 2.74 m/s . Their 75.9 kg owner has the same momentum as the two pets taken together.

Find the direction and magnitude of the owner's velocity.
Answer #1

M = 27.2 Kg
vi = 2.64 j^ m/s
m = 5.85 Kg
vi = 2.74 i^ m/s

px = 5.85 * 2.74 = 16.03 i^
py = 27.2 * 2.64 = 71.8 j^


p = sqrt(px^2 + py^2)
p = sqrt(16.03^2 + 71.8^2)
p = 73.56 kg m/s

Magnitude of owner's Velocity = 73.56/75.9
Magnitude of owner's Velocity = 0.969 m/s

Direction = tan^-1(71.8/16.03)
Direction =  77.4o